OpenStudy (anonymous):

Divide 4.48 by 2.56.

OpenStudy (anonymous):

And yes I know how to divide. :) haha This one hasn't been coming out right though and I could use some quick help.

OpenStudy (unklerhaukus):

\[\frac{4.48}{2.56}=\frac{448}{256}=\frac{7}{4}\]

OpenStudy (anonymous):

Isn't it 1.75? I have the answer but can't get it to work out quite yet.

OpenStudy (unklerhaukus):

yes 7/4 =1+3/4 is equivalent to 1.75

OpenStudy (anonymous):

What do you mean by 1+3/4?

OpenStudy (unklerhaukus):

\[\frac74=\frac{4+3}{4}=\frac44+\frac34=1+\frac34\]

OpenStudy (anonymous):

Thank you. :)

OpenStudy (unklerhaukus):

seven quarters is one and three quaters